Credit Sales     N$ 582 000

Mark-up          30%

Opening Stock    N$ 38,400

Closing Stock     N$ 18,800

Calculate the following

Gross profit                                                  

                                          



1
Expert's answer
2021-08-11T09:50:44-0400

"Cost\ of\ goods\ sold=\frac{sales}{mark-up}\\
=\frac{582000}{1.3}=447692.31\\
Gross\ profit= cost\ of\ goods\ sold\times mark\ up\\
=447692.31\times 0.3\\
= 134307.69"